OpenStudy (anonymous):

if i add 2 million to bag a then bag a would exactly double bag b if i add a and b together i get c knowing c is worth 50 million what are the values of a and b

OpenStudy (anonymous):

The unit would be millions. add 2 million to bag a then bag a would exactly double bag b aka + 2 + a = 2 b 2+a=2b if i add a and b together i get c + a + b = c or a+b=c knowing c is worth 50 million c = 50 So in summary, 2+a=2b a+b=c, which then turns into a+b=50 because you know that c=50 Find a and b.